OpenStudy (anonymous):

Solution set for (2x-5)(3x+1)=0

OpenStudy (anonymous):

Something times zero always equals something. So either (2x-5)=0 And/OR (3x+1)=0

OpenStudy (anonymous):

So start with 2x-5=0 2x-5+5=0+5 2x=5 2x/2=5/2 x=5/2

OpenStudy (anonymous):

But wait ... there's more... Do the same with 3x+1=0 This will give you the other solution